OpenStudy (anonymous):

anti-derivative of f'(x) = 0

OpenStudy (anonymous):

@beccaboo333

OpenStudy (beccaboo333):

I have no idea how to do this kind of math.. I'm sorry :c

OpenStudy (anonymous):

it's okay :)

OpenStudy (amistre64):

what derivaitve rule gets us a derivative of 0?

OpenStudy (anonymous):

it's chain right?

OpenStudy (amistre64):

in general ... no

OpenStudy (anonymous):

Could you hint me? i'm a little bad with calc.

OpenStudy (amistre64):

the hint will be: tell me the derivative rules that you rememebr

OpenStudy (anonymous):

chain rule quotient rule product rule

OpenStudy (amistre64):

exponent rule, log rule ... none of those say f'(x)=0

OpenStudy (amistre64):

the most basic rule is for something of the form: f(x) = x^n

OpenStudy (anonymous):

the part that confuses me is "f'(x)" = 0 because the antiderivative of 0 is 0 + c.

OpenStudy (amistre64):

they split this rule into two forms simply becuase they dont like x^0 for x=0, otherwise its the same

OpenStudy (amistre64):

and 0+c is just some constant

OpenStudy (anonymous):

oh I was off with that. :/

OpenStudy (amistre64):

\[y=x^0\]\[y'=0x^{-1}=0\] but if x=0 then we get a divide by 0 that everyone frowns upon so they simply make the rule that x^0 is some constant, or just 1 such that: \[y=kx^0\] \[y=k*1\] \[\frac{d}{dx}(y)=\frac{d}{dx}k*1\] \[\frac{d}{dx}(y)=k*\frac{d}{dx}1\] \[y'=k*0\]

OpenStudy (anonymous):

so the final turns out to be k * 0, not 0 + c right?

OpenStudy (amistre64):

0+c is fine, or k+c, or just C

OpenStudy (anonymous):

f'(x)= k * 0

OpenStudy (anonymous):

tks

OpenStudy (amistre64):

yep